OpenStudy (anonymous):

Which of the following is the rational exponent expression of sixth root of f?

OpenStudy (anonymous):



OpenStudy (anonymous):

As I said before... =) \[\Large a \cdot {x^{n/m}} = a \cdot \sqrt[m]{{{x^n}}}\]

OpenStudy (anonymous):

its hella confusing

OpenStudy (anonymous):

\[\Large \sqrt[6]{{{f^1}}} = {f^{1/6}}\]